OpenStudy (anonymous):

5 6 - = ------------ 3 x+2 SOLVE FOR X. NEED THE STEPS

OpenStudy (anonymous):

Do cross multiplication first.

OpenStudy (anonymous):

PLEASE HELP

OpenStudy (anonymous):

5 6 - = ------------ 3 x+2 SOLVE FOR X. NEED THE STEPS

OpenStudy (anonymous):

So, you will get the following:\[5(x+2)=6(3)\]\[5x+10=18\]Then simplify the equation.\[5x=18-10=8\]\[x={8\over5}\]

OpenStudy (anonymous):

THX